OpenStudy (anonymous):

Who ever solves these for me will get MEDAL AND FAN <3 6= a/4 + 2 -6 + x/4 = -5 9 x -7 = -7 0= 4 + n/5 -4 = r/20 -5 -1 = 5+x/6

OpenStudy (leader):

1. 4=a/4 a=16

OpenStudy (leader):

2. x/4=1 x=4

OpenStudy (leader):

3. 9x=0 x=0

OpenStudy (leader):

4. n/5=-4 n=-20

OpenStudy (leader):

5. 1=r/20 r=20

OpenStudy (leader):

6. -6=x/6 x=-36

OpenStudy (anonymous):

Thanks! :)

OpenStudy (wmj259):

Well done @Leader.

OpenStudy (leader):

No pron.

OpenStudy (leader):

Thanks

OpenStudy (leader):

prob*

OpenStudy (wmj259):

@Leader, will you help her on test day?

OpenStudy (anonymous):

what?

OpenStudy (leader):

Nah...

OpenStudy (leader):

That is cheating..

OpenStudy (anonymous):

I know but what does he mean by Test day

OpenStudy (anonymous):

Anyway thx @Leader

OpenStudy (leader):

no prob

OpenStudy (leader):

You can close this.

OpenStudy (leader):

@Kim21

OpenStudy (leader):

@Kim21 Please close this. Thanks

OpenStudy (anonymous):

sorry @Leader I had to go away for a while